HUBBLE SPACE TELESCOPE DAILY REPORT #2329 PERIOD COVERED: 0000Z (UTC) 03/15/99 - 0000Z (UTC) 03/16/99 Daily Status Report as of 075/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD/WF/PC-2 7615 (Determination of the Mass Function of a Globular Cluster Through Gravitational Microlensing) The Space Telescope Imaging Spectrograph (CCD) and the WF/PC-2 were used to make gravitational microlensing observations of the large globular cluster M-22 located in Sagittarius in order to study its mass function. Gravitational microlensing offers a unique and direct method of determining the masses of stars and other compact objects down to planetary masses. A bias and dark calibration observations were also taken. The observations were completed as planned, and no problems were reported. 1.2 Completed Two Sets of WF/PC-2 8053 (WF/PC-2 Cycle-7 Earth Flats -- Part II) The WF/PC-2 was used to take Earth flat calibration images. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Six Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the star HAR01-16 as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD 7634 (CCD Flat-Field Stability and Cosmetic Defect Fraction) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor imaging flat-field stability and cosmetic defects over monthly periods though out Cycle 7. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.8 Completed WF/PC-2 6668 (Star Formation and Galaxy Morphologies in Distant, X-ray Luminous Clusters of Galaxies) The WF/PC-2 was used to examine distant rich clusters of galaxies which are natural laboratories for studying galaxy evolution because they provide large galaxy samples at fixed redshifts. Spectacular WFPC images of distant clusters of galaxies have revealed surprisingly complex galaxy morphologies and much higher interaction rates than in present-day clusters. The obseervation completed as planned with no reported problems. 1.9 Completed Three Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.10 Completed STIS/MA1 8397 (STIS MSM Update Spectroscopic Verification) The Space Telescope Imaging Spectrograph (MA1) was used to obtain wavecals taken with G140L and G140M to check the positioning of the spectrum on the detector after the MSM update. The proposal completed as planned with no reported problems. 1.11 Completed Three Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.12 Completed STIS/CCD 8393 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation IV) The Space Telescope Imaging Spectrograph (CCD) was used to make POMS, non-scripted, low galactic latitude observations made in parallel with the prime science observations. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 7467 (Do Massive Star Clusters Form in Young and Weak Galaxy Interactions?) The WF/PC-2 was used to image two interacting galaxy pairs in two bandpasses {B and I} to determine whether the super star clusters that are so prevalent in strong, evolved interactions have also been able to form in the weak or young interactions. Secondly, if such clusters are found, it is hoped to correlate their properties as much as possible to the strengths and ages of the interactions. The proposal completed with no anomalous activity. 1.14 Completed Two Sets of WF/PC-2 6815 (The Origin of the Polarization in the Gravitationally Lensed "Cloverleaf" BAL QSO) The WF/PC-2 was used to obtain polarimetry that will explore the polarization mechanism operating in the gravitationally lensed "Cloverleaf" broad absorption line {BAL} QSO H1413+1143. The observations were completed, and there were no reported anomalous activity. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared twice. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH